OpenStudy (anonymous):

what is the slope of the line represened by the equation y=x/3-5

OpenStudy (anonymous):

y = mx + b m is the slope

OpenStudy (anonymous):

\[\huge y = \frac{ 1 }{ 3 }x - 5 \]

OpenStudy (shamil98):

Your equation is written in slope-intercept form which is: \[\huge y = mx + b\] \[\huge y = \frac {1}{3}x -5\] Where m = slope So what is your slope?

OpenStudy (abb0t):

3.1